A payday loan company charges a $60 fee for a $700 payday loan that will be repaid in 11 days.
Treating the fee as interest paid, what is the equivalent annual interest rate?
____ % interest Round to the nearest percent. This will be over 100%

Answer by Theo(13342)

60/700 = interest rate of .0857142857 per 11 days.
if you treat this as simple interest, then the equivalent annual interest will be 60 * 365 / 11 = 1990.909091, if we assume 365 days to a year.
that would be an annual interest rate of 1990.909091 / 700 * 100 = 284.4155844% per year.